OpenStudy (anonymous):

if 82degrees + 0.35x = arctan(cot 0.45x) , what is x?

hartnn (hartnn):

but how would u proceed....? not the best way, i think.

OpenStudy (ajprincess):

yup u r right. that's vat I am thnkng.

hartnn (hartnn):

ArcTan[Cot[y]] = y - Pi/2 arctan(cot 0.45x) = 0.4x -90

hartnn (hartnn):

oh, wait. tan(90-y)= cot y, right ? so arctan(cot y) should be 90-y arctan(cot 0.45x) =90-0.4x

hartnn (hartnn):

0.45x

hartnn (hartnn):

90-0.45x=82+0.35x

hartnn (hartnn):

u getting x=10 ?

OpenStudy (anonymous):

yap !! :)) thanks!

hartnn (hartnn):

messed up last problem, atleast could solve this.....

OpenStudy (anonymous):

totally fine :)) i'll answer that later on

OpenStudy (ajprincess):

U r very brilliant @hartnn

OpenStudy (anonymous):

that's right!

hartnn (hartnn):

thanks :D

OpenStudy (ajprincess):

yw nd u deserv that compliment:)